The average of two positive integers is 134. One of the numbers is the square root of the other. Find the integers.
Found 2 solutions by ganesh, stanbon:
Answer by ganesh(20)


Let the numbers be x and y and y = sqrt{x}.
Therefore, y - sqrt{x}= 0--------(1)
Also, {y + x}/2 = 134
i.e., y + x = 268---------(2)
Solving (1) and (2), we get x + sqrt{x}= 268
i.e., (x-268)^2 = {-sqrt{x})^2
so, x^2 -537+ 81796 = 0
Solving these equation, we get one positive root.
That is the value of x and using this value find the value of y.

Answer by stanbon(75887)

The average of two positive integers is 134. One of the numbers is the square root of the other. Find the integers.
-------------------
(x+sart(x))/2 = 134
x + sqrt(x) = 268
x + sqrt(x)-268 = 0
Let w = sqrt(x)
-----------------
x^2+w-268 = 0
w = [-1 +- sqrt(1-4*-268)]/2
w = [-1 +- sqrt(1073)]/2
w = [1 +- 32.756678...]/2
w = 16.87833935... or w = -15.87833935...
-----------
Comment:
These answers are not even close to being integers.
Check your text problem or your posting for errors.
==========================
Cheers,
Stan H.